# The Simons Observatory: Validation of reconstructed power spectra from simulated filtered maps for the Small Aperture Telescope survey

### v1

@@ 3.3.1 Time-domain noise simulations
  • Scan Synchronous Signal (SSS): We use an ad hoc model where the telescope scans a fixed map in horizontal coordinates to emulate both optical pickup and signals internal to the instrument. All detectors observe the same map, but the map is drawn separately for each simulated observation. The map itself is Gaussian white noise, convolved with a $1^{\circ}$ beam, inversely proportional to the observing elevation and scaled to 1 mK RMS at $45^{\circ}$ observing elevation.
@@ 4.1 Comparison of band power window functions
For the TF method, we filter our power-law simulations with the observation matrix, so we can guarantee the exact same filtering operations between both methods. We run the estimation of the TF with 100 simulations of a power-law with spectral index $-2$ and the same binning scheme.
@@
We also simulate CMB realisations, generated with the Planck best-fit cosmology [59], including $B$-modes arising only from gravitational lensing ($r=0$), as described in Section 3.1. We generate them at $N_{\rm side}=128$ and filter them with the same scheme represented by the observation matrix. The reconstructed polarisation spectra with both methods are shown in Figure 4. The four square panels show the $EE$, $EB$, $BB$, and $BE$ spectra in clockwise order. In each, we show the mean $C_{\ell}$ across $N_{\rm sims}=100$ realisations. The error bar is the error of the mean, i.e. the standard deviation across realisations divided by $\sqrt{N_{\rm sims}}$. The thinner lower panel on each spectrum shows the ratio between spectrum residuals and the associated error. The residual $\Delta(C_{\ell})$ is the difference between the measured decoupled power spectrum and the fiducial theory spectrum binned using the BPWFs. The observed CMB maps include beam smoothing, and so the theory spectra are convolved with a $30^{\prime}$ Gaussian beam.
@@ 4.2 Full multipole range using the transfer function method
The TF-deconvolved spectra for the observed CMB and for the full range $\ell=30-300$, are shown in Figure 6. For each spectrum we show the mean $C_{\ell}$ with error on the mean across $N_{\rm sims}=100$ realisations. The right-hand side panel shows the residuals divided by the error bar. No systematic bias in the residual is visible, except for the first band power $\ell=30-40$. As discussed earlier, this bias is insignificant when considering the real 1-year SAT observation.
In Section 5, we will simulate a null test, allowing us to assess the robustness of the TF method with noiseless and noisy data.
@@ 5.2 CMB-only test
If the residuals are consistent with noise, the $\chi^{2}$ statistic of eq. (5.2) should be consistent with a chi-squared distribution with 27 degrees of freedom (DOF), this being the number of band powers in the range $\ell=30-300$. We can calculate the probability-to-exceed (PTE) of each realisation. The histogram of the 100 PTEs is shown on the left-hand side of Figure 10 for the three $XY$ polarisations: $EE$, $EB$, and $BB$. If the residual in this null test is consistent with noise, we would expect a uniform distribution of PTEs between 0 and 1.
@@ 5.3 Signal and noise test
Our survey will not be pure CMB signal but also contain inhomogeneous and non-Gaussian polarised foregrounds, as well as noise. We include the pysm d1 and s1 foregrounds. The main issue of concern for our analysis is the spatial inhomogeneity and non-Gaussian nature of these templates, which could potentially induce biases in the deconvolved spectra due to a TF calculation that relies on spatially homogeneous Gaussian simulations, as described in Section 3.1. We use these models with moderate levels of SED variability as we care more about spatial than spectral complexity. We leave the detailed study of foreground-induced biases in a TF reconstruction for future work.
@@ 6 Full time-domain simulations
Since we have all the components necessary to simulate an entire SO-SAT deep survey, we could run a limited-scale test of our complete analysis, from observing data all the way to measuring a tensor-to-scalar ratio value. Since we only simulate two of the SO frequency bands, f090 and f150, we use Planck observations as ancillary data to perform component separation. [^15] The real SO deep survey will eventually observe 6 frequency bands in total and will not need to use ancillary data from other experiments. Since this test has a limited scale, we compromise by only simulating the 353 GHz Planck band. On the other hand, we do not include synchrotron as a polarised foreground in our simulation, since we do not include the low-frequency bands necessary to constrain it, but only thermal dust.
The CMB realisations we use only have lensing $B$-modes and $r=0$. The thermal dust model is d1 from pysm. In Section 3.1 we described how these simulations are generated. We observe the 353 GHz Planck simulations with the same schedule and filtering described in Sections 3.2 and 3.3 with toast. In this paper, we work under the simplified scheme that f090 and f150 have been observed and filtered in the same way. This can be easily generalised to the point where each frequency channel is assigned its own filtering recipe. This filtering scheme is applied over an entire simulated Planck 353 GHz map, which includes the noise. We have 100 realisations of the npipe 353 GHz frequency maps, with their A/B detector split. For the two SAT bands, we define four quarter-mission (QM) splits. The pure power-law simulations are co-added with these same splits, and the TFs are calculated for the corresponding split combinations and used to calculate the corresponding cross-spectra. Finally, we estimate unbiased spectra by averaging all cross-spectra combinations, avoiding all auto-spectra. For example, the f090$\times$×f090 spectrum is calculated from the 6 cross-spectra combinations between the 4 QM splits, the f090$\times$×f150 spectrum is calculated from the 16 cross-spectra combinations, etc. For spectra, we use the full South patch mask constructed from the hits map shown in Figure 2.
To estimate a covariance for error bars and parameter posteriors with Markov chain Monte Carlo (MCMC) sampling, we use 100 realisations of Gaussian dust and calculate their deconvolved spectra in the same way as described above.
@@
Our next step is to do component separation and estimate posterior probabilities for parameters, in particular $r$. We do this with the spectra-based bbpower pipeline,161616github.com/simonsobs/BBPower which we have used before to forecast the performance of SO [69, 70, 71, 72]. We fit a model only to the $BB$ spectrum, with five free parameters: $r$, a lensing amplitude $A_{\rm lens}$, a dust power spectrum amplitude (at pivot scale $\ell=80$ and pivot frequency $353$ GHz) $A_{\rm d}^{BB}$, a power-law index of the dust power spectrum $\alpha_{\rm d}^{BB}$, and a frequency spectral index $\beta_{\rm d}$, corresponding to the assumed MBB spectral energy distribution of Galactic dust emission. The temperature $T_{\rm d}=20$ K of the dust MBB is fixed in our model. For each parameter, we have a fiducial value. The CMB realisations are generated with no primordial $B$-modes ($r=0$), and $A_{\rm lens}=1$, while d1 dust has spatially variable SED spectral parameters and equation (3.1) is a good approximation within small patches on the sky, but it is not exact for a larger patch such as the South Galactic patch. We fit these parameters to the pure unfiltered d1 template within our common hits map at the three relevant frequencies: 93, 145, and 353 GHz, computing spectra in the multipole range $\ell=30-300$ and within our mask. As error bars for the fit, we use the standard deviation across the 100 realisations of unfiltered Gaussian dust. We use the curve_fit method from scipy [73] to perform the fit. These fiducial values and priors are listed in Table 2 for the five parameters. While we list the dust fiducial values with their $1\sigma$ uncertainties, we use the fitted value alone.
@@
In order to focus on the ability of our pipeline to obtain unbiased constraints from the first year of SO-SAT data, and not on the forecasted sensitivity, we present our simulated cosmological constraints normalised by their standard deviation, referred to in this work as the normalised bias. [^17] We subtract the fiducial value from the parameter and divide by the $1\sigma$ error bar. We note, nevertheless, that the obtained constraints are in broad agreement with the forecasts presented in Refs. [2, 23]. The histogram for the normalised bias for the best-fit value for each parameter is shown in Figure 12, right. Figure 13 shows the posterior distributions for our five parameters for the 100 realisations. For each parameter value, we subtract the fiducial value and divide by the $1\sigma$ uncertainty.

### v2

@@ 3.3.1 Time-domain noise simulations
  • Scan Synchronous Signal (SSS): We use an ad hoc model where the telescope scans a fixed map in horizontal coordinates to emulate both optical pickup and signals internal to the instrument (e.g. electromagnetic interference from the azimuthal drive). All detectors observe the same map, but the map is drawn separately for each simulated observation. The map itself is Gaussian white noise, convolved with a $1^{\circ}$ beam, with amplitude inversely proportional to the observing elevation and scaled to 1 mK RMS at $45^{\circ}$ observing elevation. The elevation dependence is added to reflect the fact that far-side lobe pickup intensifies at lower observing elevations. The $1^{\circ}$ smoothing is chosen to avoid sharp features in the simulated signal. Finally, the overall amplitude of the SSS is chosen to overwhelm CMB signal for the sky modes that are degenerate with it. These choices are anecdotally supported by polarbear data but are not based on rigorous analysis. The model will be thoroughly calibrated against real data in the future.
@@ 3.3.2 Observation matrix
The filtering matrix, $\bm{\mathrm{Z}}$, operating on a TOD vector, $\bm{d}$, can be understood to perform three separate operations:
  1. 1.
    Perform a generalized least squares fit of $\bm{d}$, in terms of the time domain templates that form the columns of $\bm{\mathrm{F}}$. This results in a vector of template amplitudes: $\bm{a}=\left(\bm{\mathrm{F}}^{T}\bm{\mathrm{N}}^{-1}_{Z}\bm{\mathrm{F}}\right)^{-1}\bm{\mathrm{F}}^{T}\bm{\mathrm{N}}^{-1}_{Z}\bm{d}$.
  2. 2.
    Unroll (weight and co-add) the templates, $\bm{\mathrm{F}}$, according to the fitted amplitudes, $\bm{a}$, to produce a model of the TOD: $\bm{d}_{\mathrm{model}}=\bm{\mathrm{F}}\bm{a}$.
  3. 3.
    Subtract the TOD model from the input to produce a filtered version of the input TOD: $\bm{d}_{\mathrm{filtered}}=\bm{d}-\bm{d}_{\mathrm{model}}$.
@@ 4.1 Comparison of band power window functions
For the TF method, we filter our power-law simulations with the observation matrix, so we can guarantee the exact same filtering operations between both methods. We run the estimation of the TF with 100 realisations of a power-law with spectral index $-2$ and the same binning scheme.
@@
We also simulate pure CMB realisations, generated with the Planck best-fit cosmology [69], including $B$-modes arising only from gravitational lensing ($r=0$), as described in Section 3.1. We generate them at $N_{\rm side}=128$ and filter them with the same scheme represented by the observation matrix. The reconstructed polarisation spectra with both methods are shown in Figure 4. The four square panels show the $EE$, $EB$, $BB$, and $BE$ spectra in clockwise order. In each, we show the mean $C_{\ell}$ across $N_{\rm sims}=100$ realisations. The error bar is the error of the mean, i.e. the standard deviation across realisations divided by $\sqrt{N_{\rm sims}}$, and as such, these $\sigma$ values are 10 times smaller than $\sigma$ for a single realisation. The thinner lower panel on each spectrum shows the ratio between spectrum residuals and the associated error. The residual $\Delta(C_{\ell})$ is the difference between the measured decoupled power spectrum and the fiducial theory spectrum binned using the BPWFs. The observed CMB maps include beam smoothing, and so the theory spectra are convolved with a $30^{\prime}$ Gaussian beam. Since these are pure CMB realizations, the only source of variance is cosmic variance.
@@ 4.2 Full multipole range using the transfer function method
The TF-deconvolved spectra for the observed CMB and for the full range $\ell=30-300$, are shown in Figure 6. For each spectrum we show the mean $C_{\ell}$ with error on the mean across $N_{\rm sims}=100$ realisations. The right-hand side panel shows the residuals divided by the error bar. No systematic bias in the residual is visible, except for the first band power $\ell=30-40$. In Section 5, we will show that this bias is sub-dominant and therefore not important when noise is added to these pure CMB simulations, through the use of null tests.
@@ 5.2 CMB-only test
If the residuals are consistent with noise, the $\chi^{2}$ statistic of Equation (5.2) should be consistent with a chi-squared distribution with 27 degrees of freedom (DOF), this being the number of band powers in the range $\ell=30-300$. We can calculate the probability-to-exceed (PTE) of each realisation. The histogram of the 100 PTEs is shown on the left-hand side of Figure 10 for the three $XY$ polarisations: $EE$, $EB$, and $BB$. If the residual in this null test is consistent with noise, we would expect a uniform distribution of PTEs between 0 and 1.
@@ 5.3 Signal and noise test
Our survey will not be pure CMB signal but also contain anisotropic and non-Gaussian polarised foregrounds, as well as noise. We include the pysm d1 and s1 foregrounds. The main issue of concern for our analysis is the spatial anisotropic and non-Gaussian nature of these templates, which could potentially induce biases in the deconvolved spectra due to a TF calculation that relies on spatially isotropic Gaussian simulations, as described in Section 3.1. We use these models with moderate levels of SED variability as our main interest is to assess the impact on mode coupling from the highly spatially anisotropic and non-Gaussian foregrounds. Using models with more SED complexity or even frequency decorrelation is beyond the scope of our paper. We leave the detailed study of foreground-induced biases in a TF reconstruction for future work.
@@ 6 Full time-domain simulations
Since we have all the components necessary to simulate an entire SO-SAT deep survey, we could run a limited-scale test of our complete analysis, from observing data all the way to measuring a tensor-to-scalar ratio value with a spectra-based component separation pipeline. Since we only simulate two of the SO frequency bands, f090 and f150, we use Planck observations as ancillary data to perform component separation. [^16] The real SO deep survey will eventually observe 6 frequency bands in total and will not need to use ancillary data from other experiments. Since this test has a limited scale, we compromise by only simulating the 353 GHz Planck band. On the other hand, we do not include synchrotron as a polarised foreground in our simulation, since we do not include the low-frequency bands necessary to constrain it, but only thermal dust.
The 100 CMB realisations we use only have lensing $B$-modes and $r=0$. The thermal dust model is d1 from pysm. In Section 3.1 we described how these simulations are generated. We observe the 353 GHz Planck simulations with the same schedule and filtering described in Sections 3.2 and 3.3 with toast. In this paper, we work under the simplified scheme that f090 and f150 have been observed and filtered in the same way. This can be easily generalised to the point where each frequency channel is assigned its own filtering recipe. This filtering scheme is applied over an entire simulated Planck 353 GHz map, which includes the noise. We have 100 realisations of the npipe 353 GHz frequency maps, with their A/B detector split. For the two SAT bands, we define four quarter-mission (QM) splits. The pure power-law simulations, used to estimate the transfer function, are co-added with these same splits, and the TFs are calculated for the corresponding split combinations and used to calculate the corresponding cross-spectra. Finally, we estimate unbiased spectra on the simulated observations of CMB+dust+noise by averaging all cross-spectra combinations, avoiding all auto-spectra. For example, the f090$\times$×f090 spectrum is calculated from the 6 cross-spectra combinations between the 4 QM splits, the f090$\times$×f150 spectrum is calculated from the 16 cross-spectra combinations, etc. For spectra, we use the full South patch mask constructed from the hits map shown in Figure 2.
To estimate a covariance for error bars and parameter posteriors with Markov chain Monte Carlo (MCMC) sampling, we use 100 realisations of Gaussian dust and calculate their deconvolved spectra in the same way as described above. The total covariance is estimated from the 100 realizations of CMB + noise + Gaussian dust.
@@
Our next step is to do component separation and estimate posterior probabilities for parameters, in particular $r$. We do this with the spectra-based bbpower pipeline,171717github.com/simonsobs/BBPower which we have used before to forecast the performance of SO [79, 80, 81, 30]. We choose this method for component separation, as it is the only foreground removal approach for which efficient and accurate methods exist to account for the impact of timestream- and/or map-level filtering [e.g. 50, 82]. We fit a model only to the $BB$ spectrum, with five free parameters: $r$, a lensing amplitude $A_{\rm lens}$, a dust power spectrum amplitude (at pivot scale $\ell=80$ and pivot frequency $353$ GHz) $A_{\rm d}^{BB}$, a power-law index of the dust power spectrum $\alpha_{\rm d}^{BB}$, and a frequency spectral index $\beta_{\rm d}$, corresponding to the assumed MBB spectral energy distribution of Galactic dust emission. The temperature $T_{\rm d}=20$ K of the dust MBB is fixed in our likelihood model. For each parameter, we have a fiducial value. The CMB realisations are generated with no primordial $B$-modes ($r=0$), and $A_{\rm lens}=1$, while d1 dust has spatially variable SED spectral parameters and Equation (3.1) is a good approximation within small patches on the sky, but it is not exact for a larger patch such as the South Galactic patch. We wish to determine these fiducial parameters for the d1 dust template. We fit these parameters to the pure unfiltered d1 template within our common hits map at the three relevant frequencies: 93, 145, and 353 GHz, computing spectra in the multipole range $\ell=30-300$ and within our mask. As error bars for the fit, we use the standard deviation across the 100 realisations of unfiltered Gaussian dust. We use the curve_fit method from scipy [83] to perform the fit. These fiducial values and priors are listed in Table 2 for the five parameters. While we list the dust fiducial values with their $1\sigma$ uncertainties, we use the fitted value alone.
@@
In order to focus on the ability of our pipeline to obtain unbiased constraints from the first year of SO-SAT data, and not on the forecasted sensitivity, we present our simulated cosmological constraints normalised by their standard deviation, referred to in this work as the normalised bias. [^18] We subtract the fiducial value from the parameter and divide by the $1\sigma$ error bar. The histogram for the normalised bias for the best-fit value for each parameter is shown in Figure 12, right. Figure 13 shows the posterior distributions for our five parameters for the 100 realisations. For each parameter value, we subtract the fiducial value and divide by the $1\sigma$ uncertainty.